OpenStudy (anonymous):

Simplify the complex fraction. y-1/y^2+y-6 over y-6/y+3

OpenStudy (anonymous):

Is this the problem?\[\frac{ \frac{ y-1 }{ y ^{2}+y-6 } }{ \frac{ y-6 }{ y+3 } }\]If so, a problem of this kind, where a fraction is divided by another fraction essentially, can be started by applying this rule: \[\frac{ \frac{ a }{ b } }{\frac{ c }{ d } }=(\frac{ a }{ b })(\frac{ d }{ c})\]

OpenStudy (anonymous):

Yes that is it. So (y-1)(y-6) and (y^2+y-6)(y+2).